OpenStudy (joftheworld):

I need help plotting three points from this linear equation please! 2y=3x+11

OpenStudy (owlcoffee):

Let's first order the line equation into it's general form: \[-3x+2y-11=0\] Now we will give the independent variable a value (the independent variable is always the "x") and see that values of "y" we get: \[x=0\] \[=> -3(0)+2y-11=0\] \[y=\frac{ 11 }{ 2 }\] So, one of the points that belong to the line is (0,11/2)

OpenStudy (joftheworld):

oh also I can't use any fractions or decimals ..

OpenStudy (owlcoffee):

Okay, then try giving it other values so you get a non-fractions, but fractions are representable, that thy don't show does not mean you cannot use them.

OpenStudy (joftheworld):

its an online test.. and I have 5 questions more.. can you please help me

OpenStudy (owlcoffee):

Yes that's why I'm here.

OpenStudy (joftheworld):

okay well I don't have much time until it ends so do you think you can just tell me the three points of each equations? I've already done the previous questions.. and it took quite sometime .. ill give you some medals if I can..

OpenStudy (owlcoffee):

I can give you the way to find them, you have to deduce the answers.

OpenStudy (joftheworld):

that works. but im still stuck on this problem

OpenStudy (owlcoffee):

What troubles you?

OpenStudy (joftheworld):

the same problem I posted above

OpenStudy (owlcoffee):

All you have to do is give any value to "x" and then solve it for "y" as I explained on my example. What i did was give x the value of zero, so I can find the "y" that corresponds to that value, therefore, finding a point that actually belongs in the line. You can give x the value of 2 and plot it in, and solve it as a 1st degree equation as you have learned in the past.

OpenStudy (joftheworld):

I know how to do it man.. but I just don't have enough time.. and im not quick at doing these equations ...

OpenStudy (anonymous):

2y = 3x + 11 y = 3/2x + 11/2 slope: 3/2 or 1.5 y-intercept: 11/2 or 5.5 can you graph from here?

OpenStudy (tkhunny):

Pick x = 1, 3, 5 and you'll be done.